OpenStudy (anonymous):

draw the graph of (x^2) + (y^2) <25

zepdrix (zepdrix):

Lets first look at x^2 + y^2 = 25 Do you know what the graph looks like? it's a special shape :)

OpenStudy (anonymous):

yah i tnk it might be the lower part of the circle

zepdrix (zepdrix):

Since it's not written explicitly as a function --> y=f(x), It is actually both the top and bottom of a circle! :) The lower part would look like, solved for y, the negative half \[y=\sqrt{25-x^2}\]

zepdrix (zepdrix):

x^2 + y^2 = r^2 <-- this is the form of a circle. So 25 = r^2 --> our radius is 5

zepdrix (zepdrix):

 The reason we color the inside, is because we shade the area LESS THAN the radius, so anything smaller than the radius. hope that makes sense :)